# MMO-IG: Multi-Class and Multi-Scale Object Image Generation for Remote Sensing

### v1

@@ IV-C Comparisons with Existing DGMs
methods aug Ground track Harbor Overpass Ship Stadium Storage tank Tennis court Train station Vehicle Windmill
R-CNN [20] 49.3 39.5 30.9 9.1 60.8 18.0 54.0 36.1 9.1 16.4
50.1 40.2 31.8 13.2 57.6 17.5 55.8 35.3 12.4 17.9
Faster R-CNN [5] 56.9 50.2 50.1 27.7 73.0 39.8 75.2 38.6 23.6 45.4
57.2 51.4 53.4 29.3 72.8 40.8 76.2 37.9 26.4 47.2
YOLO [21] 61.1 44.9 49.7 87.4 65.9 68.7 83.4 29.4 48.3 78.7
69.2 39.8 53.4 90.5 70.6 76.3 87.3 26.4 52.3 80.6
PANet [22] 73.4 45.3 56.9 71.7 70.4 62.0 80.9 57.0 47.2 84.5
70.2 44.9 57.3 72.8 71.7 64.2 81.2 58.3 48.2 85.3
CornerNet [23] 79.5 26.1 60.6 37.6 70.7 45.2 84.0 57.1 43.0 75.9
82.3 31.2 62.2 38.3 71.3 46.2 85.2 61.5 44.5 76.2
@@ V Conclusion
In this paper, we propose a generative framework for synthesizing RS images containing MMOs with corresponding instance-level labels to support research on RSIOD. In the framework, ISIM and SODI are introduced as control conditions for encoding the MMOs and image content from local and global perspectives simultaneously. Meanwhile, SCDKG is proposed to model the complex interdependencies among objects of different classes and their diverse spatial geometric characteristics, which ensure a rational and realistic content distribution for the generated RS images. Experiments demonstrate the effectiveness of MMO-IG for generating high-quality RS images and the ability to improve the detection performance of RSIOD by alleviating data limitation problems.

### v3

@@ IV-C Comparisons with Existing DGMs
methods aug Ground track Harbor Overpass Ship Stadium Storage tank Tennis court Train station Vehicle Windmill Avg
R-CNN [7] 49.3 39.5 30.9 9.1 60.8 18.0 54.0 36.1 9.1 16.4 37.1
50.1 40.2 31.8 13.2 57.6 17.5 55.8 35.3 12.4 17.9 37.7
Faster R-CNN [5] 56.9 50.2 50.1 27.7 73.0 39.8 75.2 38.6 23.6 45.4 53.4
57.2 51.4 53.4 29.3 72.8 40.8 76.2 37.9 26.4 47.2 54.0
YOLO [21] 61.1 44.9 49.7 87.4 65.9 68.7 83.4 29.4 48.3 78.7 55.3
69.2 39.8 53.4 90.5 70.6 76.3 87.3 26.4 52.3 80.6 55.7
PANet [23] 73.4 45.3 56.9 71.7 70.4 62.0 80.9 57.0 47.2 84.5 62.9
70.2 44.9 57.3 72.8 71.7 64.2 81.2 58.3 48.2 85.3 63.2
CornerNet [25] 79.5 26.1 60.6 37.6 70.7 45.2 84.0 57.1 43.0 75.9 62.4
82.3 31.2 62.2 38.3 71.3 46.2 85.2 61.5 44.5 76.2 62.5
@@ V Conclusion
In this paper, we present a generative framework, MMO-IG, designed to synthesize remote sensing (RS) images with multi-modal objects (MMOs) and provide corresponding instance-level labels for Remote Sensing Image Object Detection (RSIOD). We introduce ISIM and SODI as control conditions to encode MMO and image content, and propose SCDKG to model interdependencies between object classes and their spatial characteristics. Experimental results show that MMO-IG effectively generates high-quality RS images and improves detection performance.
However, MMO-IG has some limitations. When handling rare instances, the generated images may fail to accurately represent them. Function fitting sometimes oversimplifies instance attributes, and the method’s performance can degrade with large numbers of targets due to increased complexity. Additionally, while MMO-IG performs well on remote sensing images, it may struggle with natural scenes that involve more complex backgrounds, lighting variations, and occlusions. Finally, the computational requirements of the image generation process limit its application in resource-constrained environments. We aim to address these challenges in future work by improving the framework’s adaptability, reducing its computational demands, and enhancing its handling of rare instances and large target counts.